# Seele: A Unified Acceleration Framework for Real-Time Gaussian Splatting

### v2

@@ Abstract
Specifically, we propose two GPU-oriented techniques: hybrid preprocessing and contribution-aware rasterization. Hybrid preprocessing alleviates the GPU compute and memory pressure by reducing the number of irrelevant Gaussians during rendering. The key is to combine our view-dependent scene representation with online filtering. Meanwhile, contribution-aware rasterization improves the GPU utilization at the rasterization stage by prioritizing Gaussians with high contributions while reducing computations for those with low contributions. Both techniques can be seamlessly integrated into existing 3DGS pipelines with minimal fine-tuning. Collectively, our framework achieves 2.6 speedup and 32.3% model reduction while achieving superior rendering quality compared to existing methods.
![[Uncaptioned image]](x1.png)
@@ 1 Introduction
3D Gaussian Splatting (3DGS) [27] has emerged as a vital technique in many rendering-related domains, including autonomous driving [36, 35, 28, 42], augmented and virtual reality (AR/VR) [41, 14, 26, 24]. These real-time applications require high rendering performance to ensure seamless quality of services. However, the limited computational resources of mobile platforms often constrain these applications, preventing them from reaching their full potential.
For instance, Nvidia AGX Orin [6], a leading-edge compute module for modern vehicles, offers only 3.4% of the computational resources of a Nvidia A100 workstation GPU [2]. Similarly, the Snapdragon XR2 chip [10, 8], the most widely used chip for AR/VR devices, has 4.2% of the computational power compared to Nvidia RTX 4090 [1]. On Nvidia’s AGX Orin, 3DGS [27] merely achieves 20 FPS on real-world datasets [13, 30, 23], far from the real-time requirement, i.e., 90 frame-per-second (FPS) of VR and gaming [7, 9, 43]. An interesting challenge is to achieve real-time performance on these devices with only 3-4% of the compute resources compared to high-end GPUs.
To achieve this goal, we must understand the major performance bottlenecks in the current 3DGS rendering pipelines. We summarized the bottlenecks into three aspects: computational intensity, rendering inefficiency, and memory budget. This paper proposes a one-stop solution framework, Seele ‡‡‡The name is an organization that comes from a famous Japanese anime, Neon Genesis Evangelion. This organization’s mission is to accomplish the Human Instrumentality Project. Similarly, the goal of our project is to address and resolve all the performance bottlenecks in 3DGS., that addresses the inefficiencies across these three dimensions (Sec. 3). Next, we briefly discuss the inefficiencies in today’s 3DGS rendering pipeline and highlight how Seele addresses these limitations.
@@
One primary challenge in the 3DGS pipeline is its computational intensity. From an individual pixel’s perspective, rendering one pixel typically involves processing thousands of Gaussian points §§§We use “Gaussians”, “Gaussian points”, and “Gaussian ellipsoid” interchangeably: there is a one-to-one mapping between them., with each GPU thread rendering one pixel. This rendering process imposes substantial computational overhead on resource-constrained devices. Existing solutions [15, 16, 37] primarily focus on reducing the overall model size, namely the number of Gaussian points, but these methods often trade off rendering quality for performance.
In contrast, we propose hybrid preprocessing with a view-dependent scene representation that dynamically loads the relevant Gaussians to GPU memory during runtime (Sec. 4). Our data representation naturally identifies the necessary Gaussians contributing to the current viewpoint while leaving other irrelevant Gaussians untouched. Meanwhile, our method does not make any compromise between performance and accuracy. On average, this method alone outperforms the state-of-the-art rendering algorithms while achieving a 2.1 speedup.
@@
The second issue with the current 3DGS rendering pipeline is that all splatted Gaussians go through the same rasterization pipeline regardless of their contribution to the final pixels. Our experiment shows that, for each pixel, 1.5% of the Gaussians contribute to 99% of the final pixel. Therefore, we argue that this uniform treatment to all Gaussians leads to severe rendering inefficiencies, as less significant Gaussians should naturally receive fewer computational budgets. However, prior work [32, 20, 33], such as pruning or quantization techniques, still retains a uniform rendering pipeline and fails to differentiate Gaussians based on their contributions.
On the contrary, we propose contribution-aware rasterization from a pixel-centric perspective (Sec. 5). Rather than reducing computation uniformly, we dynamically identify and prioritize Gaussians with high contributions (i.e., high accumulative transparencies) and reduce computation budgets for those with low contributions. Gaussians with low contributions, which typically contribute to low-frequency textures (as shown in Fig. 4), are allocated less computation by dynamically skipping insignificant operations. We show that this method achieves an additional 1.3 speedup while retaining the same rendering quality.
@@
Our last contribution eases the peak GPU memory. As model complexity scales up, the GPU memory requirements often become unsustainable due to the increasing number of Gaussians involved. Current solutions [32, 40] apply offline compression techniques to reduce memory usage, but they overlook a key opportunity: rendering a given viewpoint does not require loading all Gaussians into GPU memory simultaneously.
Leveraging this key observation, we design proactive memory management that asynchronously loads necessary data to GPU memory without disrupting the critical path of rendering. Combined with our view-dependent scene representation, our memory management significantly reduces peak GPU memory usage. On average, our approach achieves a 32.3% reduction in model size, allowing large-scale rendering on mobile devices.
Collectively, our framework achieves 2.6 speedup and 32.3% runtime model reductions, all while retaining a better rendering quality compared to state-of-the-art algorithms. Furthermore, our techniques are orthogonal to existing optimizations and can be seamlessly integrated with other mainstream Gaussian splatting pipelines.
@@ 2.1 Efficient Data Representation
A 3DGS model is essentially a set of points in 3D space that represent the physical structure of the world. The model size often becomes one of the key bottlenecks in rendering performance and memory usage. To reduce the model size, several studies have proposed various data representations to optimize storage and performance [15, 16, 20, 32, 39]. Some reduce GPU memory usage by leveraging encoding techniques to compress the model. For example, CompactGS [32] employs vector quantization to compress the model offline and decodes it at runtime using a pretrained codebook. Similarly, EAGLES [20] uses an encoding network to compress the neural radiance field into a compact representation. However, both methods introduce non-trivial execution overhead during runtime, which can limit their applicability in computation-constrained scenarios.
Other approaches, such as LightGaussian[15], Mini-Splatting[16], and Taming3DGS [37], propose various pruning techniques to eliminate insignificant Gaussians. However, these methods have to make a trade-off between rendering quality and efficiency. Some studies adopt structured data representations, such as octrees [40] and kd-tree [28], to organize Gaussian points more effectively. However, the irregular tree traversals required by these structures introduce additional runtime overhead.
@@ 2.2 Rasterization Optimization
In addition to pruning Gaussian points, another line of research focuses on optimizing the performance of the 3DGS pipeline itself. Recent studies propose various online filtering techniques, such as axis-aligned bounding box (AABB) [29, 44] and oriented bounding box (OBB) [21, 31] intersection tests, to perform more fine-grained filtering before rasterization. These works aim to reduce the number of Gaussians processed during the rasterization stage at runtime, thereby reducing the computational workload.
On the other hand, methods like FlashGS [17] and Balanced3DGS [22] focus on optimizing the rasterization stage. For example, FlashGS designs a software pipelining to overlap data fetching with rasterization, improving overall efficiency. In contrast, Balanced3DGS addresses workload imbalance by introducing an offline scheduling mechanism to rebalance workloads at both the block and tile levels. Additionally, a few studies [31, 34, 18] explore the new hardware designs to support 3DGS, further advancing the efficiency and scalability of the pipeline.
@@ 3.1 Preliminaries
Once all Gaussians are sorted, the rasterization stage renders these Gaussians tile-by-tile. Within each tile, every pixel iterates through the same set of Gaussians, computing transparency and accumulating their colors into the pixel in the sorted order. Eqn. 1 governs the color accumulation process of pixel p:
where denotes the accumulative transmittance of pixel p from the Gaussian st to the th. and stand for the transparency and the color at the th Gaussian, respectively. The transparency of a Gaussian is calculated by the Gaussian opacity and the 2D covariance matrix [45]. is a function of the Gaussian’s spherical harmonics.
Note that, if the Gaussian’s value is insignificant (less than , i.e., ), this Gaussian will be skipped in the color accumulation to avoid numerical instabilities. The color accumulation process terminates once the accumulative transmittance surpasses a predefined threshold, .
@@ 3.2 Overview
Fig. 2 gives the overview of our acceleration framework, Seele. To address three rendering inefficiencies highlighted in Sec. 1, we introduce two key components: hybrid preprocessing and contribution-aware rasterization to reduce the runtime computation and peak memory.
In 3DGS pipelines, rasterization is the most time-consuming stage, accounting for 67% of the execution time on a Nvidia Ampere GPU [6]. Prior studies have explored two primary methods to reduce the overall workload of rasterization. The first approach applies pruning techniques, which eliminate insignificant Gaussians offline [15, 16]. The second proposes online filtering techniques, such as AABB and OBB intersection tests, to perform more fine-grained filtering before rasterization [29, 44, 21, 31].
Our hybrid preprocessing combines the merits of both offline and online techniques. During the offline processing, we design a view-dependent scene representation that clusters Gaussians based on their contributions to a set of closely related rendering viewpoints. At runtime, our framework can asynchronously prefetch the relevant clusters into the GPU memory before rendering. This approach largely reduces the filtering workload during preprocessing. In addition, the view-dependent scene representation inherently eliminates Gaussians that are irrelevant to current rendering while within the view frustum. Combined with online filtering, our hybrid preprocessing collectively minimizes the processed Gaussians in rasterization.
@@
Specifically, our rasterization pipeline organizes pixels into small groups while still keeping one thread responsible for one pixel. At runtime, only a single pixel within each pixel group calculates its Gaussian transparency rather than having all pixels compute . The purpose is to let this pixel identify Gaussians that potentially contribute less to the pixel group. Since the identified Gaussians are likely to have minimal impact on the remaining pixels within a pixel group, we exclude these Gaussians from further rasterization. In this way, we eliminate insignificant Gaussian accumulation at runtime. Combined with our integrated co-training, we can achieve even better rendering quality.
4 Hybrid Preprocessing, HP
@@
Based on this observation, our data representation classifies Gaussians into two categories: shared Gaussians and exclusive Gaussians, as shown in Fig. 3. Shared Gaussians are shared across all viewpoints and are frequently used for rendering frames from any viewpoint. On the other hand, exclusive Gaussians are specifically attached to a group of viewpoints and are only used for rendering frames within that group. For instance, Fig. 3(a) illustrates the original set of Gaussian points, which are then clustered as shown in Fig. 3(b). The yellow points in Fig. 3(b) represent the shared Gaussians, while the remaining colored points correspond to the exclusive Gaussians in different clusters.
Clustering Strategy.
The key to our approach lies in how to classify the Gaussians into shared and exclusive Gaussians. Initially, our method randomly samples camera poses and clusters them into smaller groups based on the weighted similarity of their camera positions, and view directions, . Specifically, we first normalize the camera positions, , then concatenate normalized and as a 1D dimensional vector, , where is a hyperparameter that balances the impact of position and orientation. is used to cluster all Gaussians.
For each cluster, we identify and reserve the Gaussians that are the primary contributors to pixel rendering. Specifically, for each pixel, we find the Top- Gaussians based on their contribution to the accumulative transmittance, , as the contribution metric. The rationale here is that the remaining Gaussians contribute minimally to the final pixel and can be safely discarded without compromising rendering quality. We empirically choose a value of 32, as we find that increasing further does not yield any significant improvement in rendering quality.
@@
Once these Gaussians are loaded, we incorporate this coarse-grained scene prefetching with fine-grained online filtering, e.g., AABB tests, as shown in Fig. 2. This further eliminates false-positive intersected Gaussians when performing Gaussian-tile intersections. The red crosses in Fig. 2 highlight the false-positive Gaussians. Sec. 7.3 shows the effectiveness of our hybrid preprocessing.
@@
Our hybrid preprocessing not only enhances rendering performance but also reduces peak GPU memory usage. However, loading Gaussian clusters at runtime could introduce non-trivial execution overhead and potentially cause frame stunning. To address this, we prefetch future clusters ahead of time by asynchronously loading them using a dedicated GPU stream. We find that using linearly extrapolating the camera pose is an effective method for predicting future poses in our scene prefetching. We adopt similar prediction methods as prior works [25, 19] and claim no contribution to this end. This way, our approach enables asynchronous prefetching to be overlapped with runtime frame rendering. Compared to loading all Gaussian clusters into GPU memory, our scene prefetching strategy reduces the runtime model size by 32.27%.
5 Contribution-Aware Rasterization, CR
5.1 Algorithm

### v3

@@ Abstract
Specifically, we propose two GPU-oriented techniques: hybrid preprocessing and contribution-aware rasterization. Hybrid preprocessing alleviates the GPU compute and memory pressure by reducing the number of irrelevant Gaussians during rendering. The key is to combine our view-dependent scene representation with online filtering. Meanwhile, contribution-aware rasterization improves the GPU utilization at the rasterization stage by prioritizing Gaussians with high contributions while reducing computations for those with low contributions. Both techniques can be seamlessly integrated into existing 3DGS pipelines with minimal fine-tuning. Collectively, our framework achieves up to 6.3 speedup and 39.1% model reduction while achieving superior rendering quality compared to existing methods. Our codes will be released upon publication.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) [kerbl20233d] has emerged as a vital technique in many rendering-related domains, including autonomous driving [liu2025citygaussian, liu2024citygaussianv2, kerbl2024hierarchical, tancik2022block], augmented and virtual reality (AR/VR) [rojas2023re, chen2023mobilenerf, hu2022efficientnerf, hedman2021baking]. These real-time applications require high rendering performance to ensure seamless quality of services. However, the limited computational resources of mobile platforms often constrain these applications, preventing them from reaching their full potential.
For instance, Nvidia AGX Orin, a leading-edge compute module for modern vehicles, offers only 3.4% of the computational resources of a Nvidia A100 workstation GPU. Similarly, the Snapdragon XR2 chip, the most widely used chip for AR/VR devices, has 4.2% of the computational power compared to Nvidia RTX 4090. On Nvidia’s AGX Orin, 3DGS [kerbl20233d] barely achieves 20 frames-per-second (FPS) on real-world datasets [barron2021mip, knapitsch2017, hedman2018deep], far from the real-time requirement, i.e., 90 FPS of VR [wang2023effect]. An interesting challenge is to achieve real-time on devices with only 3-4% of the resources compared to high-end GPUs.
To achieve this goal, we must understand the major performance bottlenecks in the current 3DGS rendering pipelines. We summarized the bottlenecks into three aspects: computational intensity, rendering inefficiency, and memory budget. This paper proposes a one-stop solution framework, Seele, that addresses the inefficiencies across these three dimensions (Sec. 3). Next, we briefly discuss the inefficiencies in today’s 3DGS rendering pipeline and highlight how our proposed solution addresses these limitations.
@@
One primary challenge in the 3DGS pipeline is its computational intensity. From an individual pixel’s perspective, rendering one pixel typically involves processing thousands of Gaussian points, with each GPU thread rendering one pixel. This rendering process imposes substantial computational overhead on resource-constrained devices. Prior solutions [fan2023lightgaussian, fang2024mini] primarily focus on reducing the overall model size, namely the number of Gaussian points, but these methods often trade off rendering quality for performance.
In contrast, we propose hybrid preprocessing with a view-dependent scene representation that dynamically loads the relevant Gaussians to GPU memory during runtime (Sec. 3.3). Our data representation naturally identifies the necessary Gaussians contributing to the current viewpoint while leaving other irrelevant Gaussians untouched. Meanwhile, our method does not make any compromise between performance and accuracy. On average, this method alone outperforms the state-of-the-art rendering algorithms while achieving a 2.8 speedup.
@@
The second issue with the current 3DGS rendering pipeline is that all splatted Gaussians go through the same rasterization pipeline regardless of their contribution to the final pixels. Our experiment shows that, for each pixel, 1.5% of the Gaussians contribute to 99% of the final pixel. Therefore, we argue that this uniform treatment to all Gaussians leads to severe rendering inefficiencies, as less significant Gaussians should naturally receive fewer computational budgets. However, prior work [lee2024compact, girish2025eagles, lin2024rtgs], such as pruning or quantization techniques, still retains a uniform rendering pipeline and fails to differentiate Gaussians based on their contributions.
On the contrary, we propose contribution-aware rasterization from a pixel-centric perspective (Sec. 3.4). Rather than reducing computation uniformly, we dynamically identify and prioritize Gaussians with high contributions (i.e., high accumulative transparencies) and reduce computation budgets for those with low contributions. Gaussians with low contributions, which typically contribute to low-frequency textures (as shown in Fig. 4), are allocated less computation by dynamically skipping insignificant operations. We show that this method achieves an additional 1.4 speedup while retaining the same rendering quality.
@@
Our last contribution eases the peak GPU memory. As model complexity scales up, the GPU memory requirements often become unsustainable due to the increasing number of Gaussians involved. Solutions [lee2024compact, ren2024octree] apply offline compression techniques to reduce memory usage, but they overlook a key opportunity: rendering a given viewpoint does not require loading all Gaussians into GPU memory simultaneously.
Leveraging this key observation, we design proactive memory management that asynchronously loads necessary data to GPU memory without disrupting the critical path of rendering. Combined with our view-dependent scene representation, our memory management significantly reduces peak GPU memory usage. On average, our approach achieves a 39.1% reduction in model size, allowing large-scale rendering on mobile devices.
Collectively, our framework achieves up to 6.3 speedup and 39.1% runtime model reductions, all while retaining a better rendering quality compared to state-of-the-art algorithms. Furthermore, our techniques are orthogonal to existing optimizations and can be seamlessly integrated with other mainstream Gaussian splatting pipelines.
@@ 2.1 Efficient Data Representation
A 3DGS model is essentially a set of points in 3D space that represent the physical structure of the world. The model size often becomes one of the key bottlenecks in rendering performance and memory usage. To reduce the model size, several studies have proposed various data representations to optimize storage and performance. Some [niedermayr2024compressed, chen2024hac, chen2025hac++, papantonakis2024reducing, durvasula2025contrags, niemeyer2024radsplat] reduce GPU memory usage by leveraging encoding techniques to compress the model. For example, CompactGS [lee2024compact] employs vector quantization to compress the model offline and decodes it at runtime using a pretrained codebook. Similarly, EAGLES [girish2025eagles] uses an encoding network to compress the neural radiance field into a compact representation. However, both methods introduce non-trivial execution overhead during runtime, which limits their applicability in computation-constrained scenarios.
Other pruning-based approaches use a learnable mask [liu2025maskgaussian, zhang2024lp, taghipour2025svr] or a significance score [fang2024mini, fang2024mini2, mallick2024taming, fan2023lightgaussian, hanson2025pup] to eliminate redundant Gaussians. However, these methods have to make a trade-off between rendering quality and efficiency. Some studies adopt structured data representations such as octrees [ren2024octree] and kd-tree [kerbl2024hierarchical] to organize Gaussian points while some researches [hamdi2024ges, held20253d, li20253d, ye2025gaussian] explore alternative primitives for more efficient scene representation. However, they introduce additional runtime overhead in loading or rasterizing Gaussians.
@@ 2.2 Rasterization Optimization
In addition to pruning Gaussian points, another line of research focuses on optimizing the performance of the 3DGS pipeline itself. Recent studies propose various online filtering techniques, such as axis-aligned bounding box (AABB) [klosowski1998efficient, wang2024adr, hanson2025speedy] and oriented bounding box (OBB) [gottschalk1996obbtree, lee2024gscore] intersection tests, to perform more fine-grained filtering before rasterization. These works aim to reduce the number of Gaussians processed during the rasterization stage at runtime, thereby reducing the computational workload.
On the other hand, some methods [feng2025flashgs, liao2025tc, gui2024balanced] focus on optimizing the rasterization stage. For example, FlashGS [feng2025flashgs] designs a software pipelining to overlap data fetching with rasterization, improving overall efficiency. In contrast, Balanced3DGS [gui2024balanced] addresses workload imbalance by introducing an offline scheduling mechanism to rebalance workloads at both the block and tile levels. Additionally, a few studies [lee2024gscore, lin2025metasapiens, pei2025gcc, lee2025vr] explore the new hardware designs to support 3DGS, further advancing the efficiency and scalability of the pipeline.
@@ 3.1 Preliminaries
Once all Gaussians are sorted, the rasterization stage renders these Gaussians tile-by-tile. Within each tile, every pixel iterates through the same set of Gaussians, computing transparency and accumulating their colors into the pixel in the sorted order. Eqn. 1 governs the color accumulation process of pixel p,
where denotes the accumulative transmittance of pixel p from the Gaussian st to the th. and stand for the transparency and the color at the th Gaussian, respectively. The transparency of a Gaussian is calculated by the Gaussian opacity , the 2D position and the 2D covariance matrix, [zwicker2001ewa],
is a function of the Gaussian spherical harmonics.
Note that, if the Gaussian’s value is insignificant (less than , i.e., ), this Gaussian will be skipped in the color accumulation to avoid numerical instabilities. The color accumulation process terminates once the cumulative transmittance surpasses a predefined threshold, .
@@ 3.2 Overview
Fig. 2 gives the overview of our proposed acceleration framework. To address three rendering inefficiencies highlighted in Sec. 1, we introduce two key components: hybrid preprocessing and contribution-aware rasterization to reduce the runtime computation and peak memory.
In 3DGS pipelines, rasterization is the most time-consuming stage, accounting for 67% of the execution time on a Nvidia Ampere GPU. Prior studies have explored two primary methods to reduce the overall workload of rasterization. The first approach applies pruning techniques, which eliminate insignificant Gaussians offline [fan2023lightgaussian, fang2024mini]. The second proposes online filtering techniques, such as AABB and OBB intersection tests, to perform more fine-grained filtering before rasterization [klosowski1998efficient, wang2024adr, gottschalk1996obbtree, lee2024gscore].
Our hybrid preprocessing combines the merits of both offline and online techniques. During the offline processing, we design a view-dependent scene representation that clusters Gaussians based on their contributions to a set of closely related rendering viewpoints. At runtime, our framework can asynchronously prefetch the relevant clusters into the GPU memory before rendering. Our view-dependent scene representation inherently eliminates Gaussians that are irrelevant to the current rendering while within the view frustum. Combined with online filtering, we further reduce the number of Gaussians intersected for each tile in the subsequent stages.
@@
Specifically, our rasterization pipeline organizes pixels into small groups while still keeping one thread responsible for one pixel. At runtime, only a single pixel within each pixel group calculates its Gaussian transparency rather than having all pixels compute . The purpose is to let this pixel identify Gaussians that potentially contribute less to the pixel group, so we can exclude these Gaussians from further rasterization. In this way, we eliminate insignificant Gaussian accumulation at runtime. Combined with our integrated co-training, we can achieve a smoother rendering experience. A more rigorous explanation of this approach is provided in Sec. 3.4.
3.3 Hybrid Preprocessing, HP
@@
Based on this observation, our data representation classifies Gaussians into two categories: shared Gaussians and exclusive Gaussians, as shown in Fig. 3. Shared Gaussians are shared across all viewpoints and are frequently used for rendering frames from any viewpoint (yellow points in Fig. 3(b)). On the other hand, exclusive Gaussians are specifically attached to a group of viewpoints and are only used for rendering frames within that group (other colors in Fig. 3(b)).
Offline Clustering.
The key to our approach lies in how to classify the Gaussians into shared and exclusive Gaussians. Initially, our method randomly samples camera poses and clusters them into smaller groups based on the weighted similarity of their camera positions, and view directions, . Specifically, we first normalize the camera positions, , then concatenate normalized and as a 1D dimensional vector, , where is a hyperparameter that balances the impact of position and orientation. Here, is set to 1 for clustering all Gaussians.
For each cluster, we identify and reserve the Gaussians that are the primary contributors to pixel rendering. Specifically, for each pixel, we find the Top- Gaussians based on their contribution to the cumulative transmittance, , as the contribution metric. The rationale here is that the remaining Gaussians contribute minimally to the final pixel and can be safely discarded without compromising rendering quality. We choose a value of 32, as we find that increasing further yields minimal quality improvement.
@@
Online Filtering.
Once these Gaussians are loaded, they are transformed into screen space with a new 2D Gaussian position and covariance matrix . To filter unnecessary Gaussians for each tile, the intersection test typically uses envelope to approximate,
where is the pixel position. Due to the redundancy of this approximation [hanson2025speedy, pei2025gcc], we further include the opacity into our online filtering equation as follows. As we would skip insignificant , which value is less than , , Eqn. 2 can be expressed as,
Therefore, we execute fine-grained online filtering by including opacity into consideration,
This further eliminates false-positive intersected Gaussians when performing Gaussian-tile intersections.
@@
Our hybrid preprocessing not only enhances rendering performance but also reduces peak GPU memory usage. However, loading Gaussian clusters at runtime could introduce non-trivial execution overhead and potentially cause frame stuttering. To address this, we prefetch future clusters ahead of time by asynchronously loading them using a dedicated GPU stream. We adopt similar prediction methods as prior works [hou2020motion, feng2024cicero] and find that linearly extrapolating the camera pose is an effective method for predicting future poses in our scene prefetching. This way, our approach enables asynchronous prefetching to be overlapped with runtime frame rendering. Compared to loading all Gaussian clusters into GPU memory, our scene prefetching reduces the runtime model size by 39.1% with a minimal runtime overhead (6% of the total latency).
3.4 Contribution-Aware Rasterization, CR